# Beyond prompt brittleness: Evaluating the reliability and consistency of political worldviews in LLMs

### v2

@@ Abstract
Due to the widespread use of large language models (LLMs) in ubiquitous systems, we need to understand whether they embed a specific “worldview” and what these views reflect. Recent studies report that, prompted with political questionnaires, LLMs show left-liberal leanings Feng et al. (2023); Motoki et al. (2024). However, it is as yet unclear whether these leanings are reliable (robust to prompt variations) and whether the leaning is consistent across policies and political leaning. We propose a series of tests which assess the reliability and consistency of LLMs’ stances on political statements based on a dataset of voting-advice questionnaires collected from seven EU countries and annotated for policy domains. We study LLMs ranging in size from 7B to 70B parameters and find that their reliability increases with parameter count. Larger models show overall stronger alignment with left-leaning parties but differ among policy programs: They evince a (left-wing) positive stance towards environment protection, social welfare state and liberal society but also (right-wing) law and order, with no consistent preferences in foreign policy and migration.
@@ 1 Introduction
Ensuring space for a plurality of ideas and opinions is crucial to keep democracies healthily operating (McQuail, 1993; Loecherbach et al., 2020). Given that one such space — the interaction with chatbots powered by LLMs — is becoming increasingly prevalent, the importance of unveiling biases in LLMs’ output is clear. In this context, people can easily ask chatbots to perform a variety of tasks such as generating arguments, summarizing and writing an email. Investigating how heterogeneous the worldviews embedded in these LLMs is, therefore, of extreme relevance for several downstream tasks, not only chatbot interactions. Continuing previous work on identifying biases in general NLP resources and models (Hovy and Prabhumoye, 2021), a number of studies have found biases of numerous types in LLMs, relating to gender Kotek et al. (2023), race (Omiye et al., 2023), culture Arora et al. (2023); Wang et al. (2023b), and political position Feng et al. (2023). Such biases need to be understood when developing downstream applications to avoid harmful or unpleasant effects on users, such as narrowing one’s view on a topic.
@@
In this paper, we focus on a specific type of bias, namely political bias in LLMs. Recent studies found that LLMs show a bias in that they agree more with left-wing political positions (Feng et al., 2023; Motoki et al., 2024). However, the scope and interpretation of these findings is not clear yet: Political positioning is an inherently multidimensional phenomenon, and while political individuals and organizations (e.g., parties) typically exhibit substantial (even if typically imperfect) internal consistency (Moskowitz and Jenkins, 2004; Tavits, 2007), this is not necessarily true for LLMs, which have only a weak notion of consistency Basmov et al. (2024). This opens two possible interpretations (and a sliding scale between them) for political bias in LLMs, which it would arguably be instructive to distinguish: (a), LLM show preferences primarily at the level of individual statements without major consideration to other, related statements; this would amount to stance taking (Küçük and Can, 2020) with respect to individual targets. Alternatively, the studies can be taken to indicate that (b), LLMs exhibit political belief systems with a substantial amount of internal consistency, similar to human political individuals – in a word, a political “worldview”, a term which is used for this concept in political science Ecker et al. (2021) and has also been suggested to apply to LLMs (Bender et al., 2021).
@@
We suggest that interpretations (a) and (b) can be distinguished with the help of two diagnostics that we take to be constituted of “worldviews”. First, the models should show high reliability in assessing political statements [^1] We adopt the term “reliability”, as consistency over testing replications, from psychometry (American Educational Research Association et al., 1999). — that is, they should give consistent answers irrespective of the formulation of the prompts. If this is not the case, models merely react to linguistic peculiarities, namely lexical choice, token or position biases (see Section 2 for details). Second, the models should show consistency in their political position, that is, should exhibit a consistent stance towards broad policy domains, with limited variance among statements within these domains, as well as a consistent commitment to a right or left leaning across domains.
@@
    We release ProbVAA, a dataset with statements on policy measures from seven EU countries with the answers (stance) from several political parties annotated on policy domains. We expand ProbVAA with paraphrased, negated, and semantically inverted versions of the core statements (§4).
@@
    We evaluate a range of SOTA LLMs on the ProbVAA dataset, finding substantial differences among LLMs with regard to reliability (§6). When evaluating stance on reliable statements (§7), we find that LLMs align more with left-leaning parties overall, but lack consistency regarding leanings: they tend to have no preference for some domains (migration, foreign policy) but agree with policies as divergent as pro-environment and law and order.
@@ 2 Related Work
The characterization of political positions is an important topic in political science, and a considerable number of computational models has shown that positions can be inferred from political texts (e.g., Laver et al., 2003; Slapin and Proksch, 2008; Glavaš et al., 2017). Comparing the positioning of political parties at low dimensional level under pre-defined scales remains an elusive goal in political science Heywood (2021). One of the most widely used scales is left-right, arguably distinguishing between progressive position (left), conservative positions (right) and compromise positions (center). Despite concerns about its validity (Kitschelt, 1994; Jahn, 2023), the scale has been validated broadly across countries (Evans et al., 1996; Budge et al., 2001) and also formed the basis for previous analyses of political bias in LLMs. An alternative to positioning actors on a global scale is to carry out a fine-grained analysis at the level of individual policy domains (Iversen, 1994; Ceron et al., 2023). For our consistency analysis in Section 7, we look at both of these levels (left-right scale and positioning within policy domains).
@@
Recent work has examined LLMs’ political ideology using global surveys such as Political Compass (Feng et al., 2023; Motoki et al., 2024; Rutinowski et al., 2023), or more country-specific questionnaires such as Pew Research’s ATP, World Values Survey (Durmus et al., 2022), and voting advice applications (VAAs) (Hartmann et al., 2023).
@@
The assessment of output variability and the quantification of model reliability in recent studies have involved the application of psychometric methods from social psychology. These studies have utilized standardized methodologies Dayanik et al. (2022) and questionnaires to create controlled environments for extracting reliable “attitudes” from LLMs (Tjuatja et al., 2023; Dominguez-Olmedo et al., 2023; Shu et al., 2023). Such approaches have proven to be instrumental in examining various societal biases in LLMs (Arora et al., 2023; Wang et al., 2023b; Hada et al., 2023; Esiobu et al., 2023; Shu et al., 2023). However, the exploration of psychometric methods to investigate political bias remains limited.
@@
There is a growing awareness of the fact that the input to an LLM is not the only, and maybe not even the most important, factor determining the output. For example, Min et al. (2022) show that swapping out gold labels for random ones only slightly reduces performance – a pattern that remains stable across almost all tested models regardless of the prompt template used. Khashabi et al. (2022) observe that soft prompts manage to solve a task even when presented as an arbitrary text, staying surprisingly close (within a 2% range) to the best soft prompt of the same size designed for that specific task. Finally, the meaning of prompts can be overshadowed by the choice of target words Webson and Pavlick (2022) which goes hand-in-hand with observed high result variance caused by recency and common token bias phenomena when the model chooses the most frequent token Zhao et al. (2021), or position bias when the model prioritizes labels that appear at a specific position Zheng et al. (2023).
@@ 3 Reliability-Aware Bias Analysis
Following up on this motivation, we now present our framework for evaluating the political bias of LLMs which involves two key elements: (1) enrichment of the dataset with prompt variations and policy-domain annotations and (2) evaluation of the reliability of answers and the consistency of stances.
@@
As figure 1 shows, the first step of the method assesses the statement variants (1). In addition to that, the reliability with respect to variations of prompt templates is evaluated by (2) using two types of instructions (personal and impersonal questions), (3) using synonyms for the response alternatives that the model should select, and (4) swapping the order of the alternatives (§ 5.2).
@@
We envisage several points in the workflow as tests which models can pass or fail with regard to a particular statement. As illustrated in figure 2, the test types are (1) robustness to sampling (with a fixed prompt), (2) robustness to paraphrasing/negation/semantic inversion of the original statement, and (3) robustness to label-order inversion in the prompt template. Only statements on which the models pass all tests are used to assess the models’ attitudes. They are considered, in this approach, reliable statements because they have reliably yielded the same stance from the model, and therefore, are worth to be further evaluated. Policy-domain annotations on the dataset make it possible to make the analysis of the reliable statements more fine-grained (§ 6 and 7).  [^2] We make the augmented dataset, including all tests, the models’ responses and code, available here https://github.com/tceron/eval_political_worldviews.
Figure 2: Overview of reliability tests
@@ 4.1 Sources
To assess the potential political worldviews embedded in LLMs, we collect a set of statements derived from Voting Advice Applications (VAAs). VAAs are tools that provide voters with insights on which parties are best aligned with their own opinions regarding policy issues. Unlike the frequently used Political Compass questionnaire, which categorizes political attitudes into a two-axis system (left/right and authoritarian/libertarian), VAAs offer a nuanced approach that ground political leanings in stances towards practical policies Palfrey and Poole (1987); Tavits (2007). These stances allow for a direct comparison of responses with those from national parties and/or candidates. On the one hand, this offers a more unbiased basis for measuring political leanings, as it does not rely on the questionnaire designer’s external classification to determine if an answer aligns with the "left" or "right" side of the political spectrum. On the other hand, it covers a wide range of policy domains that varies from environmental protection to government expenditures, providing more fine-grained insights on the types of biases.
@@
4.2 Policy-Domain Annotation
@@
Table 1: Fleiss $\kappa$ between three annotators for policy-domain annotations.
We have enriched ProbVAA with policy-domain annotations based on the pattern of the Swiss VAA, SmartVote [^3] More info on https://www.smartvote.ch/en/wiki/methodology-smartspider/23_ch_nr?locale=en_CH. It contains annotations that allow for the visualization and deeper understanding of the positioning of parties according to predominant policy-domains in the political spectrum. We draw from the documentation provided by SmartVote where eight categories (considered stances on policy domains) are defined: open foreign policy, liberal economic policy, restrictive financial policy, law and order, restrictive migration policy, expanded environmental protection, expanded social welfare state, and liberal society. These categories are based on domains identified in Swiss political spectrum Hermann and Leuthold (2001, 2003), but that are generalized across European countries, as evidenced by the similarity with domains analyzed in cross-European studies such as the Chapel Hill Survey Jolly et al. (2022).
When answering ‘agree’ to a statement emphasizes any of the eight given policies, the statement is marked as a ‘agree’ with that policy domain, while disagreements with a policy are annotated as ‘disagree’. Three annotators with background in traditional or computational political science extended the annotations to the other countries. Table 1 shows that inter-annotator agreement – which is calculated with agreement between ‘agree’, ‘agree’ and no label per statement – is good. The final gold annotations are drawn from the majority votes. Note that some statements do not fall into any category. Therefore, the gold annotations contain 193 statements in total (tables 9 and 10, Appendix A provide examples and details).
@@ 5.1 Models
Given that we formulate our prompts as zero-shot instruction, we opt for the instruction-following model families that vary in parameter size and contain different sets of pretraining data  [^4] We also evaluated the base models, but they do not produce any reliable answers, and are therefore not reported here.. Specifically, we focus on all size variants of LLama-2-Chat (7B, 13B, 70B) Touvron et al. (2023), the XXL variant of Flan-T5 (11B, Chung et al., 2022), Mistral Instruct (7B, Jiang et al., 2023) for open-sourced models and GPT-3.5 as a closed-source model. The models form three natural classes: small ($<$10B parameters, mistral-7b and llama2-7b), medium (between 10 and 19B (flanT5-xxl-11b and llama2-13b) and big ($>$20B, gpt3.5-20b and llama2-70b). All models utilize the top-p nucleus decoding approach, suggested by Holtzman et al. (2020), when generating responses, aiming to capture the model’s stance distribution through the sampling of the output (Cf. Appendix B for information on the implementation).
@@ 5.2 Prompt Design
For clarity, we define a prompt as composed of two parts, the prompt template (which contains the instruction given to the model) and the statement (an instance from ProbVAA).
When designing model inputs, we aim at creating templates that reliably elicit responses from models that can be clearly aligned with a defined stance, so either positive or negative. [^5] An example of an invalid response is I don’t know or I don’t have personal opinions. Considering recent research findings indicating that the meaning of prompts can be overshadowed by the choice of the target words as discussed in § 2, we aim at diversifying the choice for prompt templates from various angles. We distinguish between personal and impersonal templates (cf. prompt variants as shown in Figure 1). Personal templates formulate the instruction as a direct question (e.g. Do you agree or disagree?, Are you pro or con?), whereas impersonal are framed as objective tasks (Analyze the following statement into the labels “favorable” or “detrimental”, Classify the following statement as…). The context for evaluating the prompts is specified as Consider the long-term societal impact…. Additionally, we vary the wording of the stance (e.g. favorable, detrimental, advantageous, disadvantageous, support, oppose) to explore potential model biases in responding to specific wordings (cf. semantic label order, Figure 1). After a pilot experiment to test which prompts elicit most valid responses, we selected 3 personal and 3 impersonal prompt templates among 8 impersonal and 6 personal templates (Appendix B.1 details the selection process). Refer to the implemented prompt templates in Table 7, Appendix A.

### v3

@@ Abstract
Due to the widespread use of large language models (LLMs), we need to understand whether they embed a specific “worldview” and what these views reflect. Recent studies report that, prompted with political questionnaires, LLMs show left-liberal leanings Feng et al. (2023); Motoki et al. (2024). However, it is as yet unclear whether these leanings are reliable (robust to prompt variations) and whether the leaning is consistent across policies and political leaning. We propose a series of tests which assess the reliability and consistency of LLMs’ stances on political statements based on a dataset of voting-advice questionnaires collected from seven EU countries and annotated for policy issues. We study LLMs ranging in size from 7B to 70B parameters and find that their reliability increases with parameter count. Larger models show overall stronger alignment with left-leaning parties but differ among policy programs: They show a (left-wing) positive stance towards environment protection, social welfare state and liberal society but also (right-wing) law and order, with no consistent preferences in the areas of foreign policy and migration.
@@ 1 Introduction
It is crucial for a democratic system to guarantee space for a plurality of ideas and opinions in all kinds of communication situations, be they political, professional or personal (Balkin, 2017). Over the last few years, one particular communication situation – interactions between chatbots powered by LLMs and their users – has become a commonplace setup for many everyday communication tasks, such as assessing arguments, summarizing texts, or writing emails Wolf and Maier (2024). Our understanding of the extent to which such LLM-based scenarios guarantee space for ideas and opinions of various kinds or, conversely, to what extent they are biased Blodgett et al. (2020), is still unfolding. Continuing work on identifying biases in previous NLP resources and models (Hovy and Prabhumoye, 2021), studies have found biases of numerous types in LLMs, including gender Kotek et al. (2023), race (Omiye et al., 2023), culture Arora et al. (2023); Wang et al. (2023b), and political position Feng et al. (2023). Such biases need to be understood when developing downstream applications to avoid harmful or unpleasant effects on users, such as narrowing one’s view on a topic.
@@
In this paper, we focus on political bias in LLMs. Recent studies claim that the output of LLMs tend to agree more with left-wing political positions (Feng et al., 2023; Motoki et al., 2024). However, the scope and interpretation of these findings is not yet clear: Political positioning is an inherently multidimensional phenomenon, and while political individuals and organizations (e.g., parties) typically exhibit substantial (even if typically imperfect) internal consistency (Moskowitz and Jenkins, 2004; Tavits, 2007), this is not necessarily true for LLMs, which have only a weak notion of consistency Basmov et al. (2024).
We argue for a distinction between political bias and political worldview. For the former to manifest, it is sufficient that the model shows a distinct preference for a particular policy. This amounts to independent stance taking (Küçük and Can, 2020) with respect to individual target statements. Arguably, this behavior constitutes a form of representation bias (Mehrabi et al., 2021; Suresh and Guttag, 2021), because when the model exhibits a preference, it reflects only one worldview rather than that of a representative sample of the population. The latter, in addition, requires consistency across a set of such policies. This is similar to how political science describes the positioning of human actors in the overall political space spanning multiple policy issues using the term “worldview” Ecker et al. (2021). The term has also been suggested to apply to LLMs (Bender et al., 2021).
@@
These characterizations suggest that political bias and political worldview can be distinguished with the help of two criteria. If an LLM fits the first, it shows political bias. If it shows both, it shows a political worldview. The first criterion is whether the models show high reliability in assessing political statements [^1] We adopt the term “reliability”, as consistency over testing replications, from psychometry (American Educational Research Association et al., 1999). – that is, whether they give consistent answers irrespective of the formulation of the prompts. If this is not the case, models merely react to linguistic peculiarities, namely lexical choice, token or (textual) position biases (see Section 2 for details). The second criterion is whether models show consistency in their political worldview: Whether they exhibit a consistent stance towards broad policy issues, with limited variance among statements within these issues or a consistent commitment to a right or left leaning across issues.
@@
    We build ProbVAA, a dataset with statements on policy measures from seven EU countries with the answers from political parties. ProbVAA contains paraphrased, negated, and semantically inverted versions of the statements, and policy issue annotations (§4).
@@
    We evaluate a range of SOTA LLMs on the ProbVAA dataset, finding substantial differences among LLMs with regard to reliability (§6). When evaluating stance on reliable statements (§7), we find that LLMs align more with left-leaning parties overall, but lack consistency regarding leanings: they tend to have no preference for some issues (migration, foreign policy) but agree with policies as divergent as pro-environment and law and order.
@@ 2 Related Work
The characterization of political positions is an important topic in political science, and a considerable number of computational models has shown that positions can be inferred from political texts (e.g., Laver et al., 2003; Slapin and Proksch, 2008; Glavaš et al., 2017). Comparing the positioning of political parties at low dimensional level under pre-defined scales remains an elusive goal in political science Heywood (2021). One of the most widely used scales is left-right, arguably distinguishing between progressive position (left), conservative positions (right) and compromise positions (center). Despite concerns about its validity (Kitschelt, 1994; Jahn, 2023), the scale has been validated broadly across countries (Evans et al., 1996; Budge et al., 2001) and also formed the basis for previous analyses of political bias in LLMs (Feng et al., 2023). An alternative to positioning actors on a scale is to carry out a fine-grained analysis at the level of individual policy issues (Iversen, 1994; Ceron et al., 2023). For our consistency analysis in Section 7, we look at both of these levels (left-right scale and positioning within policy issues).
@@
Recent work has examined LLMs’ political ideology using surveys such as Political Compass (Feng et al., 2023; Motoki et al., 2024; Rutinowski et al., 2024), or more country-specific questionnaires such as Pew Research’s ATP, World Values Survey (Santurkar et al., 2023), and voting advice applications (VAAs) (Hartmann et al., 2023).
@@
The assessment of output variability and the quantification of model reliability in recent studies have involved the application of psychometric methods from social psychology. These studies have utilized standardized methodologies Dayanik et al. (2022) and questionnaires to create controlled environments for extracting reliable “attitudes” from LLMs (Tjuatja et al., 2023; Dominguez-Olmedo et al., 2023; Shu et al., 2024). Such approaches have proven to be instrumental in examining various societal biases in LLMs (Arora et al., 2023; Wang et al., 2023b; Hada et al., 2023; Esiobu et al., 2023; Shu et al., 2024). However, the exploration of psychometric methods to investigate political bias remains limited.
@@
There is a series of studies suggesting that the input to an LLM plays an important role in determining its output. For example, Min et al. (2022) show that swapping out gold labels for random ones only slightly reduces performance – a pattern that remains stable across almost all tested models regardless of the prompt instruction used. Khashabi et al. (2022) observe that continuous prompts manage to solve a task even when presented as an arbitrary instruction, staying surprisingly close (within a 2% range) to the best prompt of the same size designed for that specific task. Finally, the meaning of prompts can be overshadowed by the choice of target words Webson and Pavlick (2022) which goes hand-in-hand with observed high result variance caused by recency and common token bias phenomena when the model chooses the most frequent token Zhao et al. (2021), or position bias when the model prioritizes labels that appear at a specific position Zheng et al. (2023).
@@ 3 Reliability-Aware Bias Analysis
Following up on this motivation, we now present our framework for evaluating the political bias of LLMs which involves two key elements: (1) enrichment of the dataset with prompt variations and policy issue annotations and (2) evaluation of the reliability of answers in terms of stances.
@@
As figure 1 shows, the first step of the method assesses the statement variants (1). In addition to that, the reliability with respect to variations of prompt instructions is evaluated by (2) using two types of instructions (personal and impersonal questions), (3) using synonyms for the response alternatives that the model should select, and (4) swapping the order of the alternatives (§ 5.2).
@@
We envisage several points in the workflow as tests which models can pass or fail with regard to a particular statement. As illustrated in figure 2, the test types are (1) robustness to sampling (with a fixed prompt), (2) robustness to paraphrasing/negation/semantic inversion of the original statement, and (3) robustness to label-order inversion in the prompt instruction. Only statements on which the models pass all tests are used to assess the models’ attitudes. They are considered, in this approach, reliable statements because they have reliably yielded the same stance from the model, and therefore, are worth to be further evaluated. policy issue annotations on the dataset make it possible to make the analysis of the reliable statements more fine-grained (§ 6 and 7).  [^2] We make the augmented dataset, including all tests, the models’ responses and code, available here: https://github.com/tceron/eval_political_worldviews.
Figure 2: Overview of reliability tests.
@@ 4.1 Sources
To assess the potential political worldviews embedded in LLMs, we collect a set of statements derived from Voting Advice Applications (VAAs). VAAs are tools that provide voters with insights on which parties are best aligned with their own opinions regarding policy issues. Unlike the frequently used Political Compass questionnaire, which categorizes political attitudes into a two-axis system (left/right and authoritarian/libertarian), VAAs offer a nuanced approach that ground political leanings in stances towards practical policies Palfrey and Poole (1987); Tavits (2007). These stances allow for a direct comparison of responses with those from national parties and/or candidates. On the one hand, this offers a more unbiased basis for measuring political leanings, as it does not rely on the questionnaire designer’s external classification to determine if an answer aligns with the “left” or “right” side of the political spectrum. On the other hand, it covers a wide range of policy issues that varies from environmental protection to government expenditures, providing more fine-grained insights on the types of biases.
@@
4.2 Policy Issue Annotation
@@
Table 1: Fleiss $\kappa$ between three annotators for policy issue annotations.
We have enriched ProbVAA with policy issue annotations based on the pattern of the Swiss VAA, SmartVote [^3] More info on https://www.smartvote.ch/en/wiki/methodology-smartspider/23_ch_nr?locale=en_CH. It contains annotations that allow for the visualization and deeper understanding of the positioning of parties according to predominant policy issues in the political spectrum. We draw from the documentation provided by SmartVote where eight categories (considered stances on policy issues) are defined: open foreign policy, liberal economic policy, restrictive financial policy, law and order, restrictive migration policy, expanded environmental protection, expanded social welfare state, and liberal society. These categories are based on policy issues identified in the Swiss political spectrum Hermann and Leuthold (2001, 2003), but that are generalized across European countries, as evidenced by the similarity with issues analyzed in cross-European studies such as the Chapel Hill Survey Jolly et al. (2022).
When answering ‘agree’ to a statement emphasizes any of the eight given policies, the statement is marked as a ‘agree’ with that policy issue, while disagreements with a policy are annotated as ‘disagree’. Three annotators with background in traditional or computational political science extended the annotations to the other countries. Table 1 shows that inter-annotator agreement – which is calculated with agreement between ‘agree’, ‘disagree’ and ‘no label’ per statement – is good. The final gold annotations are drawn from the majority votes. Note that some statements do not fall into any category. Therefore, the gold annotations contain 193 statements in total (tables 9 and 10, Appendix A provide examples and details).
@@ 5.1 Models
Given that we formulate our prompts as zero-shot instruction, we opt for the instruction-following model families that vary in parameter size and contain different sets of pretraining data  [^4] We also evaluated the base models, but they do not produce any reliable answers, and are therefore not reported here.. Specifically, we focus on all size variants of LLama-2-Chat (7B, 13B, 70B) Touvron et al. (2023), the XXL variant of Flan-T5 (11B, Chung et al., 2024), Mistral Instruct (7B, Jiang et al., 2023) for open-sourced models and GPT-3.5 as a closed-source model. The models form three natural classes: small ($<$10B parameters, mistral-7b and llama2-7b), medium (between 10 and 19B (flanT5-xxl-11b and llama2-13b) and big ($>$20B, gpt3.5-20b and llama2-70b). All models utilize the top-p nucleus decoding approach, suggested by Holtzman et al. (2020), when generating responses, aiming to capture the model’s stance distribution through the sampling of the output (Cf. Appendix B for information on the implementation).
@@ 5.2 Prompt Design
For clarity, we define a prompt as composed of two parts, the prompt instruction (which contains the instruction given to the model) and the statement (an instance from ProbVAA).
When designing model inputs, we aim at creating templates that reliably elicit responses from models that can be clearly aligned with a defined stance, so either positive or negative. [^5] An example of an invalid response is I don’t know or I don’t have personal opinions. Considering recent research findings indicating that the meaning of prompts can be overshadowed by the choice of the target words as discussed in § 2, we aim at diversifying the choice for prompt instructions from various angles. We distinguish between personal and impersonal templates (cf. prompt variants as shown in Figure 1). Personal templates formulate the instruction as a direct question (e.g. Do you agree or disagree?, Are you pro or con?), whereas impersonal are framed as objective tasks (Analyze the following statement into the labels “favorable” or “detrimental”, Classify the following statement as…). The context for evaluating the prompts is specified as Consider the long-term societal impact… Additionally, we vary the wording of the stance (e.g. favorable, detrimental, advantageous, disadvantageous, support, oppose) to explore potential model biases in responding to specific wordings (cf. semantic label order, Figure 1). After a pilot experiment to test which prompts elicit most valid responses, we selected 3 personal and 3 impersonal prompt instructions among 8 impersonal and 6 personal templates (Appendix B.1 details the selection process). Refer to the implemented prompt instructions in Table 7, Appendix A.